# Transportability of Prognostic Markers: Rethinking Common Practices through a Sufficient-Component-Cause Perspective

### v2

@@ Abstract
Transportability, the ability to maintain performance across populations, is a desirable property of of markers of clinical outcomes. However, empirical findings indicate that markers often exhibit varying performances across populations. For prognostic markers that are advertised as predictive risk equations, oftentimes a form of updating is required when the equation is transported to populations with different disease prevalences. Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC). We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity). Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions. A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods. For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation. Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on the distribution of causes. A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
footnotetext: From Faculty of Pharmaceutical Sciences and Faculty of Medicine (MS), Vancouver, Canada; School of Population Health, WHO Collaborating Centre for Climate Change and Health Impact Assessment, and enAble Institute, Faculty of Health Sciences, Curtin University, Perth, Australia (GP); Faculty of Medicine, Universitas Negeri Malang, Indonesia (GP); Saw Swee Hock School of Public Health, National University of Singapore, Singapore (WC)footnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, British Columbia, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
@@
The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are transportable from one population to another. In this context, we define transportability as conditions under which certain indices of marker performance remain the same across populations. Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration. However, in practice, this premise more often than not fails to hold. Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs1. For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)2. The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population. A similar phenomenon was observed in risk prediction models for acute exacerbations of obstructive lung disease3. In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
@@
Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as case-mix- is available. One of the most common information pieces is the outcome prevalence. For example, cancer registries often provide a good estimate of cancer risk in a population. This poses the challenge as to how such information can be used to revise our assessment of marker performance when transporting it to a new population. For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5. When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7. The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that are distributed, and act, independently of predictors used in the model6. Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
@@
The transportability of predictive information across populations is an active area of research in predictive analytics and machine learning. A recent scoping review categorized methods aimed at developing transportable marker, or making an existing marker transportable to a new setting, based on whether they require access to data from the target population, and whether they are purely data-driven or require contextual knowledge about associations9. One common underlying framework in knowledge-driven approaches is causal graphs, particularly the pioneering work on selection diagrams by Pearl et al.10. Generally, these methods aim to identify and remove predictors whose association with the outcome varies across populations. More recent developments include the creation of graph surgery estimators that break the link between selection into the population and causal associations between the outcome and predictors whose association with the outcome is not consistent11,12. This approach has recently been used to study how common metrics of model performance for prognostic and diagnostic markers change across populations with difference case-mix13.
@@
Causal graphs are not the only model for causation. Another is the sufficient component causes (SCC) model14,15. SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs). This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
@@
In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome. As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue. We later discuss how symmetries can be used to investigate the transportability of diagnostic markers but leave deeper investigations to future work. This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence. Our thesis is that by reducing the transportability problem to its basic constituents in this model, patterns will emerge that can provide insight into more complex scenarios.
@@
The SCC framework assumes the existence of sets of sufficient causes that bring about an event18. Within each set, the causes are non-redundant (all elements with the set are required for the event to happen), but sets can act independently of each other19.
Consider a binary prognostic marker, such as the presence or absence of BRCA1 mutation, and a binary outcome, such as the occurrence of breast cancer. While BRCA1 mutation is associated with increased cancer risk20, it is not a definitive marker: neither a negative BRCA1 mutation eliminates the risk of breast cancer, nor does its presence guarantee that cancer will occur. The fact that neither positive nor negative marker values are definitive indicates that there are at least two other mechanisms at play. On one hand, a BRCA1-positive individual must experience some other, key events that ultimately lead to cancer, explaining why not everyone with a BRCA1 mutation will develop cancer. On the other hand, cancer can also occur via pathways independent of the BRCA1 mutation, explaining why some breast cancer cases are BRCA1-negative.
We now formalize a minimal setup that explains such indeterminism of a marker value for an outcome. We are interested in the properties of a binary prognostic marker $T$ for a binary outcome $D$. For this marker to be informative but not definitive, at least two latent variables (or switches) must be present that can cause false-negative and false-positive responses. We can model these switches as follows:
@@
Our interest is in the four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$). By convention PPV and NPV are referred to as predictive values and SE and SP as accuracy metrics.
Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$. However, this is mathematically symmetrical to our reference setup. This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$. Because of this, any pattern we observe for PPV in the main formulation is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa. Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup. Table 1 shows the minimum SCC setup, and the resulting 2-by-2 contingency tables of marker and outcome values ($P(T,D)$), from which performance metrics can be derived.
@@
If all we know about marker performance in a populations is its contingency table, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three. However, any joint distribution can be replaced by independent marginal distributions for $P(T=1)$, $P(U=1)$, and $P(V=1)$ that would result in the same contingency table. We use the shorthand notations $P_{T}=P(T=1),P_{U}=P(U=1)$, and $P_{V}=P(V=1)$. Given specifying three Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table. This 1:1 mapping can thus be formulated as follows:
@@
In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV). Ultimately, using a marker for decision-making entails interpreting its predictive values - i.e., $P(D|T)$. Different transportation methods construct these predictive values by combining different pieces of information from the source and target populations.
@@
Transportation by predictive values refers to the where predictive values from the source population are considered to be the same as in the target population. Given that $P(T,D)=P(T)P(D|T)$, this transportability means our estimate of the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population. This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D=1|T)$ for someone with marker value $x$. For a binary marker, this risk equation can be written as P(D=1|T)=(1NPV)+[PPV(1NPV)]TP(D=1|T)=(1-{NPV})+[{PPV}-(1-{NPV})]T. In our numerical example, advertising prognostic information as a risk equation would result in $P(D=1|T)=0.375+0.375T$.
@@
Thus, under this transportability scheme we need access to prevalence in the target population to derive predictive values. This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22. Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
Returning to our running example, imagine we are transporting the above-mentioned prognostic marker to a population where $P_{T}=$ 0.250, $P_{U}=$ 0.750, $P_{V}=$ 0.500. The true contingency table and marker performance metrics are provided in Table 2. Transportation by predictive values preserves PPV and NPV in the target population, but will result in SE and SP of 0.5 and 0.833. Now, imagine we know the target population prevalence, which is 0.622 (32.7% higher than in the source population). If we consider transportation by accuracy, we will arrive at the PPV and NPV values of, respectively, 0.750 and 0.625.
Prevalence adjustment for risk equations
A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models. A familiar modeling framework for binary outcomes is the logistic regression. There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability. Applying this function to both sides of the previous equation, we have
@@
The last term on the right-hand side is the likelihood ratio (LR) of the marker between the diseased and non-diseased groups. The first term on the right-hand side is the logit of prevalence, which is not a function of marker value. These derivations indicate that the practice of prevalence adjustment by the odds-ratio of prevalence between the source and target populations is equivalent to the application of Bayes’ theorem (where outcome prevalence is taken as the prior probability). This approach changes the locus of transportability from predictive values to the LR: for prevalence adjustment to make the marker transportable, the LR must remain constant across all levels of the marker.
For binary markers, the LR is defined at two values. For $T=0$ it is (1SE)/SP(1-SE)/SP (aka negative LR), and for $T=1$ it is SE/(1SP)SE/(1-SP) (aka positive LR). For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant. This in turn implies that prevalence adjustment for binary markers is equal to transportation by accuracy; i.e., considering SE and SP to be transportable.
@@
The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23. While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic. Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations. One can indeed create population-generating mechanisms where a given subset of these metrics remains constant while others vary. However, the SCC framework helps us examine the plausibility of such mechanisms.
Table 3 shows conditions under which each of the four metrics remains constant (and therefore transportable) for a prognostic marker.
Table 3: Conditions for Transportability of Metrics of Marker Performance
@@
  • As an example of derivations, consider PPV. Its transportability means $P(D=1|T=1)=c$ (a constant). But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)$.
For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant. For NPV, transportability requires that among test-negative individuals, the proportion in whom both the universal and alternative causes are present should remain constant. These conditions are both satisfied if the prevalences of $U$ and $V$ are stable across populations. In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$. In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate. For many outcomes, including breast cancer, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24. In comparison, conditions for the transportation by accuracy are more complicated. It does not seem possible to specify a simple model on how $P(T,U,V)$ should vary such that both SE and SP remain stable. Such transportability method for a prognostic marker, as is implied in prevalence adjustment, amounts to a very specific set of assumptions on how causes vary across populations.
@@
In the SCC framework, the degree of transportability of a marker by a given performance metric depends on how the causes vary across populations. Despite the simple setup, complex patterns arise due to the non-linearities in the interplay among causes. In this section we visualize how the relationship between performance metrics and outcome prevalence changes under various population-generating mechanisms. Taking the population of our numerical example as the baseline, we modeled the following scenarios: when populations vary only in one of causes; when they vary in two of the three causes; and when they vary in all three causes. When multiple causes vary, we assume they change by the same degree on the odds-ratio scale.
Results are provided in Figure 1. As explained above, under $T$-only variation, PPV and NPV remain transportable. This again highlights that transporting a marker by its predictive values (as in risk equations) means attributing change in prevalence entirely due to $T$, a very strong assumption in most real scenarios. Importantly, in all other conditions, PPV and NPV varied by prevalence. On the other hand, in none of the modeled scenarios did SE and SP remain unchanged. In fact, SE and SP could vary in either direction as a function of prevalence, reflecting the more complex conditions required for their transportability.
@@
The SSC framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions. Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations. In this case, it is reasonable to assume that $P_{T}$ remains essentially constant across populations. When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence. The resulting transportation method will be different from transportation by predictive values or accuracy. Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
@@
But, what if the only information we have is differences in outcome prevalence across populations? We believe it is still worth visiting the plausibility of the assumptions under various transportability methods even under such a general case. For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence. Transportability by accuracy is also questionable due to unclear assumptions it places on the distribution of causes. For this general case, it might be more reasonable to take a neutral yet transparent stance about the variations in causes, for example that all causes move by the same extent to cause change in prevalence. As an implementation of this approach, instead of applying a correcting odds-ratio to predicted values, one can solve for a common odds-ratio that, when applied to $P_{T}$, $P_{U}$, and $P_{V}$ in the source population, results in a new population that matches the outcome prevalence in the target population. This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods. Details of operationalizing this approach are provided in Appendix 2.
In our numerical example (last row of Table 2), we need to apply an odds-ratio of 1.612 to the three causes to match the target population’s outcome prevalence. This odd-ratio can be derived via standard univariate root finding - given that there is a monotonical relationship between this odds-ratio and prevalence. It can also be expressed as the real root of a cubic equation for which standard solutions are available (see Appendix 2). The contingency table under this method of transportability will be {0.290, 0.060, 0.333, 0.318}. This in turn results in SE of 0.465 and SP of 0.841. The PPV and NPV are, respectively, 0.829 and 0.489. Formulating marker performance as a risk equation would thus result in $P(D|T)=0.171+0.317x$. Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
@@
Ultimately, any assumption about how causes vary across populations will be a simplified version of a complex reality. The discrepancy between marker performance under a given transportability assumption and the true marker performance results in loss of prognostic information. We conducted brief simulation studies to explore such information loss under simple population-generating scenarios. The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error). A more foundational approach is to measure information loss according to information theory. In this context, the relevant measure would be the Kullback-Leibler divergence ($D_{KL}$)25. This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26. In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.

### v3

@@ Abstract
Transportability, the ability to maintain performance across populations, is a desirable property of markers of clinical outcomes. However, empirical findings indicate that markers often exhibit varying performances across populations. For prognostic markers that are advertised as predictive risk equations, oftentimes a form of updating is required when the equation is transported to populations with different disease prevalences. Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC). We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity). Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions. A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods. For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation. Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on the distribution of causes. A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
Keywords: Prognosis; Predictive Value of Tests; Sensitivity and Specificity; Biomarkers; Etiology; Epidemiologic Methods
footnotetext: From Faculty of Pharmaceutical Sciences and Faculty of Medicine, Vancouver, Canada (MS); School of Population Health, WHO Collaborating Centre for Climate Change and Health Impact Assessment, and enAble Institute, Faculty of Health Sciences, Curtin University, Perth, Australia (GP); Faculty of Medicine, Universitas Negeri Malang, Indonesia (GP); Saw Swee Hock School of Public Health, National University of Singapore, Singapore (WC)footnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, British Columbia, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
@@
The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are transportable from one population to another. However, in practice, this premise more often than not fails to hold. Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs4. For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)5. The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population. In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model6.
@@
When transporting a marker for an outcome to a new population, some information about the characteristics of that population - generally referred to as case-mix- is often available. One of the most common information pieces is the outcome prevalence. For example, cancer registries often provide a good estimate of cancer risk in a population. This poses the question as to how such information can be used to revise our assessment of marker performance. For risk predictions models that return a quantitative estimate of outcome risk given patient characteristics, some form of model revision is often needed to correct for the under-estimation or over-estimation of risks in the new population7. When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,8,9. Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any risk prediction algorithm, including black-box (e.g., machine learning) models10.
@@
The transportability of predictive information across populations is an active area of research in predictive analytics and machine learning. A recent scoping review categorized methods aimed at developing transportable marker, or making an existing marker transportable to a new setting, based on whether they require access to data from the target population, and whether they are purely data-driven or require contextual knowledge about associations11. One common underlying framework in knowledge-driven approaches is causal graphs, particularly the pioneering work on selection diagrams by Pearl et al.12. Generally, these methods aim to identify and remove predictors whose association with the outcome varies across populations13,14. Causal diagrams have recently been used to study how common metrics of model performance for prognostic and diagnostic markers change across populations with difference case-mix15.
@@
Causal graphs are not the only model for causation. Another is the sufficient component causes (SCC) model16,17. SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs). This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association18 and more biology-aligned statistical models19.
@@
In this paper, we use a parsimonious SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome. As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue. This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence. Our thesis is that by reducing the transportability problem to its basic constituents in this model, patterns will emerge that can provide insight into more complex scenarios.
@@
The SCC framework assumes the existence of sets of sufficient causes that bring about an event20. Within each set, the causes are non-redundant (all elements with the set are required for the event to happen), but sets can act independently of each other22.
Consider a binary prognostic marker, such as the presence or absence of BRCA1 mutation, and a binary outcome, such as the occurrence of breast cancer. While BRCA1 mutation is associated with increased cancer risk23, it is not a definitive marker: neither a negative BRCA1 mutation eliminates the risk of breast cancer, nor does its presence guarantee that cancer will occur. The fact that neither positive nor negative marker values are definitive indicates that there are at least two other mechanisms at play. On one hand, a BRCA1-positive individual must experience some other, key events that ultimately lead to cancer, explaining why not everyone with a BRCA1 mutation will develop cancer. On the other hand, cancer can also occur via pathways independent of the BRCA1 mutation, explaining why some breast cancer cases are BRCA1-negative.
We now formalize a minimal causal setup. We consider a binary prognostic marker $T$ for a binary outcome $D$. For this marker to be informative but not definitive, at least two latent variables (or switches) must be present that can cause false-negative and false-positive responses. We model these switches as follows:
@@
Our interest is in the transportability of four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$). By convention PPV and NPV are referred to as predictive values, and SE and SP as accuracy metrics.
Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$. However, this is mathematically symmetrical to our reference setup, as the complementary marker (whose positive and negative results are swapped) in this setup is a marker for not experiencing the outcome: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$. Because of this, any pattern we observe for PPV in the main formulation is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa. A similar symmetry exists between SE and SP in these setups. Table 1 shows these minimum SCC setups and the resulting contingency tables ($P(T,D)$).
@@
If all we know about marker performance in a populations is its contingency table, the full joint distribution of these causes among individuals in that population is unidentifiable, given its specification requires seven degrees of freedom, but the contingency table only offers three. However, any joint distribution can be replaced by independent marginal distributions for $P(T)$, $P(U)$, and $P(V)$ that would result in the same contingency table. Therefore, for studying contingency tables (and resulting marker performance metrics) for a single population, it is sufficient to model the probability of these causes independently across individuals within that population. However, the prevalence of these causes across populations can be correlated, and it is this dependence structure that determines the transportability of performance metrics.
Within a population, given specifying three independent Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table. We use the shorthand notations $P_{T}=P(T=1),P_{U}=P(U=1)$, and $P_{V}=P(V=1)$. The 1:1 mapping can be formulated as follows:
@@
Ultimately, using a marker for decision-making entails interpreting its predictive values - i.e., $P(D|T)$. Different transportation methods construct these predictive values by combining different pieces of information from the source and target populations.
@@
Transportation by predictive values refers to the where predictive values from the source population are considered to be the same as in the target population. Given that $P(T,D)=P(T)P(D|T)$, this transportability means the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population. This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D=1|T)$ for someone with marker value $x$. For a binary marker, this risk equation can be written as P(D=1|T)=(1NPV)+[PPV(1NPV)]TP(D=1|T)=(1-{NPV})+[{PPV}-(1-{NPV})]T. In our numerical example, advertising prognostic information as a risk equation would result in $P(D=1|T)=0.375+0.375T$.
@@
Thus, under this transportability scheme we need access to prevalence in the target population to derive predictive values. This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome24,25. Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
Returning to our running example, imagine we are transporting the above-mentioned prognostic marker to a population where $P_{T}=$ 0.250, $P_{U}=$ 0.750, $P_{V}=$ 0.500. Outcome prevalence in this population is 0.622 - 32.7% higher than in the source population. Transportation by predictive values preserves PPV and NPV in the target population, but will result in SE and SP of 0.5 and 0.833. Now, imagine we know the target prevalence. If we consider transportation by accuracy, we will arrive at the PPV and NPV values of, respectively, 0.750 and 0.625. The contingency table and marker performance metrics for the true population and those implied by the above-mentioned transportability methods are provided in Table 2.
Prevalence-adjustment for risk equations
A common way that prognostic markers are transported is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models. A familiar modeling framework for binary outcomes is the logistic regression. There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability. Applying this function to both sides of the previous equation, we have
@@
The last term on the right-hand side is the likelihood ratio (LR) of the marker between the diseased and non-diseased groups. The first term on the right-hand side is the logit of prevalence, which is not a function of marker value. These derivations indicate that the practice of prevalence-adjustment by the odds-ratio of prevalence between the source and target populations is equivalent to the application of Bayes’ theorem. This approach changes the locus of transportability from predictive values to the LR. For binary markers, the LR is defined at two values. For $T=0$ it is (1SE)/SP(1-SE)/SP (aka negative LR), and for $T=1$ it is SE/(1SP)SE/(1-SP) (aka positive LR). For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant.
@@
The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence26. While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic. Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations. One can create population-generating mechanisms where a given subset of these metrics remains constant while others vary. However, our parsimonious framework helps us examine the plausibility of such mechanisms.
Table 3 shows conditions for the distribution of causes across populations under which each of the four metrics remains constant (and therefore transportable) for a prognostic marker.
Table 3: Conditions for Transportability of Metrics of Marker Performance (c indicates a constant value)
@@
  • As an example of derivations, consider PPV. Its transportability means $P(D=1|T=1)=c$ (a constant). But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)=P(U=1)$ (the last equality is based on our assumption of independence of cause distributions within a population).
For the PPV to be transportable, the prevalence of universal causes should remain constant across populations. For NPV, transportability requires that the proportion in whom both the universal and alternative causes are present should remain constant. These conditions are both satisfied if $P_{U}$ and $P_{V}$ are stable across populations. In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$. In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate. For many outcomes, this assumption is unrealistic (e.g., many factors contribute to the variability in breast cancer risk27). In comparison, conditions for the transportation by accuracy are more complicated. For SE, it requires the ratio of the probability of one cause ($V$) over the odds of another ($T$) to remain constant. For SP, it does not seem possible to specify a simple model on how causes should vary to ensure transportability. Overall, transportation by accuracy metrics for prognostic markers amounts to placing a very specific set of conditions on how causes vary across populations.
@@
In our simple causal framework, the degree of transportability of a marker by a given performance metric depends on how the causes vary across populations. Despite the simple setup, complex patterns arise due to the non-linearities in the interplay among causes. In this section we visualize how the relationship between performance metrics and outcome prevalence changes under various population-generating mechanisms. Taking the population of our numerical example as the baseline, we modeled the following scenarios: when populations vary only in one of causes; when they vary in two of the three causes; and when they vary in all three causes. When multiple causes vary, we assume they change by the same degree on the odds-ratio scale.
Results are provided in Figure 1. As explained above, under $T$-only variation, PPV and NPV remain transportable. Importantly, in all other conditions, PPV and NPV varied by prevalence. On the other hand, in none of the modeled scenarios did SE and SP remain unchanged. In fact, SE and SP could vary in either direction as a function of prevalence, reflecting the more complex conditions required for their transportability.
@@
This framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions. Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations. In this case, it is reasonable to assume that $P_{T}$ remains essentially constant across populations. When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence. The resulting transportation method will be different from transportation by predictive values or accuracy metrics. Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
@@
Methods based on causal graphs provide specific solutions for marker transportability if causal associations between specific factors that determine the marker and outcome values are known. But, what if the only information we have is differences in outcome prevalence across populations? Our parsimonious framework enables an overall assessment of the plausibility of the assumptions under various transportability methods even under such a general case. For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence. Transportability by accuracy is also questionable due to unclear assumptions it places on the distribution of causes. For this general case, it might be more reasonable to take a neutral yet transparent stance about the variations in causes, for example that all causes move by the same extent to cause change in prevalence. As an implementation of this approach, instead of applying a correcting odds-ratio to predicted values, one can solve for a common odds-ratio that, when applied to $P_{T}$, $P_{U}$, and $P_{V}$ in the source population, results in a new population that matches the outcome prevalence in the target population. This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods. Details of solving for this common odds-ratio are provided in the Appendix.
In our numerical example (last row of Table 2), we need to apply an odds-ratio of 1.612 to the three causes to match the target population’s outcome prevalence. The contingency table under this method of transportability will be {0.290, 0.060, 0.333, 0.318}. This in turn results in SE of 0.465 and SP of 0.841. The PPV and NPV are, respectively, 0.829 and 0.489.
@@
Ultimately, any assumption about how causes vary across populations will be a simplified version of a complex reality. The discrepancy between marker performance under a given transportability assumption and the true marker performance results in loss of prognostic information. We conducted brief simulation studies to explore such information loss under simple population-generating scenarios. The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on certain aspects of performance (e.g., discrimination, calibration, or prediction error). A more foundational approach is to measure information loss by the Kullback-Leibler divergence ($D_{KL}$), an information-theoretic measure of the discrepancy between a true distribution and a candidate distribution28. $D_{KL}$ quantifies the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself29. In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.